OpenStudy (cutiecomittee123):

Solve for x logx+log(x-1)-log(4x)

OpenStudy (joannablackwelder):

Hm, is there an equals sign in there somewhere?

OpenStudy (cutiecomittee123):

Ooops!! logx+log(x-1)=log(4x)

OpenStudy (cutiecomittee123):

@JoannaBlackwelder

OpenStudy (joannablackwelder):

Ah, that makes more sense. :-) Any ideas how to start?

OpenStudy (cutiecomittee123):

Uhm....no lol

OpenStudy (joannablackwelder):

Ok, no worries. We need to combine the logs on the left side.

OpenStudy (joannablackwelder):

loga + logb = log(ab)

OpenStudy (joannablackwelder):

Can you use that formula?

OpenStudy (cutiecomittee123):

Got it so logx multiplies by log(x-1) which is log x(x-1) right?

OpenStudy (joannablackwelder):

Yep, great!

OpenStudy (joannablackwelder):

So, we have log x(x-1) = log (4x)

OpenStudy (joannablackwelder):

Any thoughts from here?

OpenStudy (cutiecomittee123):

Maybe multiply the left side

OpenStudy (cutiecomittee123):

to get log (x^2-1)

OpenStudy (joannablackwelder):

Multiplying sounds good. But check your distributing.

OpenStudy (cutiecomittee123):

Wait so then would it be x^2-x??

OpenStudy (cutiecomittee123):

log x^2-x?

OpenStudy (joannablackwelder):

Yep :-)

OpenStudy (joannablackwelder):

Ok, so, any ideas on the next step?

OpenStudy (cutiecomittee123):

divide by 4 on the left side?

OpenStudy (joannablackwelder):

Hm, no, we can't divide. Because we still have logs.

OpenStudy (cutiecomittee123):

Okay well...

OpenStudy (joannablackwelder):